OpenStudy (zairahernandez):

find three points for the linear equation -x+2y=11?

OpenStudy (owlcoffee):

You can find any point belonging to any geometric place, as long as you have the equation and it is set to the general form. What we will do is assign values to "x" and solve for "y" in order to obtain a point that belongs to the line, in this case: \[-x+2y=11\] taking it to it's general form: \[-x+2y-11=0\] So I will assign a value to "x", let's say... x=1: \[-(1)+2y-11=0\] \[-1+2y-11=0\] \[2y=12\] \[y=\frac{ 12 }{ 2 }\] \[y=6\] So, concluding, one of the points that beling to the line, let's call it point "P" is: P(1,6)

OpenStudy (zairahernandez):

Thank YOU